What is the molarity of OH− in solutions
with an H3O+ concentration of 0.00037 mol/L?
Answer in units of mol/L
1
Expert's answer
2020-05-24T16:31:56-0400

Concentration of "H_3O^+=0.00037\frac{mol}{L}"

"pH=-log(0.00037)=3.43"

"pOH=14-3.43=10.57"

Hence, molarity of "OH^-=antilog(-10.57)=0.269\times10^{-10}\frac{mol}{L}"